OpenStudy (anonymous):

Simplify the expression. Write the answer using scientific notation. (9 x 10^7) (7 x 10^9)

OpenStudy (anonymous):

You always use the same method for this kind of problem. Step one: regroup your multiplicators (9 and 7 in your case) Step two: regroup your powers of 10 (10^7 and 10^9) so, you get: 9*7=63 10^7*10^9=10^16 (simply add the exponents) In scientifical notation, we don't want numbers such as 63... it's too big! But, we know that 63=6,3*10^1, so, we repeat the same process we just did : 6,3*10^1*10^16=6,3*10^17 Is this clear?

OpenStudy (anonymous):

yes(: